The control panel on a spaceship contains a light that blinks every 2.00 s as observed by an astronaut in the ship. If the spaceship is moving past Earth with a speed of 0.750 c, determine the time interval between blinks as observed by a person on Earth.


c = 3times10^8


1
Expert's answer
2022-04-22T16:43:21-0400

"t=\frac{t_0}{\sqrt{1-\frac{v^2}{c^2}}}=3~s."